# ALI-DPFL: Differentially Private Federated Learning with Adaptive Local Iterations

### 25 Feb 2024 v4

@@ Abstract
We consider differentially private federated learning in a resource-constrained scenario, where both privacy budget and communication rounds are constrained. By theoretically analyzing the convergence, we can find the optimal number of local DPSGD iterations for clients between any two sequential global updates. Based on this, we design an algorithm of Differentially Private Federated Learning with Adaptive Local Iterations (ALI-DPFL). We experiment our algorithm on the MNIST, FashionMNIST and Cifar10 datasets, and demonstrate significantly better performances than previous work in the resource-constraint scenario. Code is available at https://anonymous.4open.science/r/ALI-DPFL-5347.
@@ I Introduction
In federated learning, each client normally downloads the global model from the center sever, performs local iterations, and uploads the resulted training parameters back to the center sever. The center sever then updates the global model accordingly. The above steps repeat until the global model converges [mcmahan2017communication]. In this way, federated learning drives by communicating only training parameters other than raw data. Nevertheless, some studies have shown that federated learning still carries privacy risks. Training parameters, such as gradient values, can be used to recover a portion of the original data [zhu2019deep] or infer whether specific content originates from certain data contributors [song2017machine]. Melis et al. [LucaMelis2022ExploitingUF] also demonstrated that participants’ training data could be leaked by shared models. Therefore, additional measures need to be taken to protect data privacy in federated learning.
Many works[AnttiKoskela2018LearningRA, LeTrieuPhong2017PrivacyPreservingDL, jayaraman2019evaluating, LiyaoXiang2019DifferentiallyPrivateDL] have demonstrated that the technique of differential privacy (DP) could protect machine learning models from unintentional information leakage. In differentially private federated learning (DPFL), each client executes a certain number of local stochastic gradient decent with differential privacy (DPSGD) iterations before performing a global aggregation. For each global aggregation, DPFL consumes privacy budget according to the number of local iterations and communicates one round of training parameters, and thus DPFL should be constraint to both privacy budget and communication rounds. However, current DPFL schemes typically optimize the global model only with the constraint of a limited privacy budget, while overlooking the constraint of communication rounds.
@@ II-B Adaptive local iterations in FL
There are currently numerous adaptive federated learning algorithms aimed at improving model performance in heterogeneous data settings. For example adaptive model aggregation [Sunwoo2023LayerWise, jayaram2022adaptive], adaptive optimizers [reddi2020FedOPT], adaptive Network Layering [karimi2021layer] and adaptive client sampling [Boxin2023AddressingBudget], etc. Wang et al. And [WangS2019ResourceConstrained] improved the federated learning algorithm by considering scenarios where clients and the server consume the same type of resources. They achieved faster convergence of the global model under resource constraints by dynamically adjusting local iterations.

### 24 Mar 2024 v5

@@ Abstract
We consider differentially private federated learning in a resource-constrained scenario, where both privacy budget and communication rounds are constrained. By theoretically analyzing the convergence, we can find the optimal number of local Differential Privacy Stochastic Gradient Descent (DPSGD) iterations for clients between any two sequential global updates. Based on this, we design an algorithm of Differentially Private Federated Learning with Adaptive Local Iterations (ALI-DPFL). We experiment our algorithm on the MNIST, FashionMNIST and Cifar10 datasets, and demonstrate significantly better performances than previous work in the resource-constraint scenario. Code is available at https://github.com/KnightWan/ALI-DPFL.
@@ I Introduction
In federated learning, each client normally downloads the global model from the center sever, performs local iterations, and uploads the resulted training parameters back to the center sever. The center sever then updates the global model accordingly. The above steps repeat until the global model converges [mcmahan2017communication]. In this way, the global model reaches convergence by communicating only training parameters instead of raw data. Nevertheless, some studies have shown that federated learning still carries privacy risks. Training parameters, such as gradient values, can be used to recover a portion of the original data [zhu2019deep] or infer whether specific content originates from certain data contributors [song2017machine]. Melis et al. [LucaMelis2022ExploitingUF] also demonstrated that participants’ training data could be leaked by shared models. Therefore, additional measures need to be taken to protect data privacy in federated learning.
Many works[AnttiKoskela2018LearningRA, LeTrieuPhong2017PrivacyPreservingDL, jayaraman2019evaluating, LiyaoXiang2019DifferentiallyPrivateDL] have demonstrated that the technique of differential privacy (DP) could protect machine learning models from unintentional information leakage. In differentially private federated learning (DPFL), each client executes a certain number of local stochastic gradient descent with differential privacy (DPSGD) iterations before performing a global aggregation. For each global aggregation, DPFL consumes privacy budget according to the number of local iterations and communicates one round of training parameters, and thus DPFL should be constrained by both the privacy budget and communication rounds. However, current DPFL schemes typically optimize the global model only with the constraint of a limited privacy budget, while overlooking the constraint of communication rounds.
@@ II-B Adaptive local iterations in FL
There are currently numerous adaptive federated learning algorithms aimed at improving model performance in heterogeneous data settings. For example adaptive model aggregation [Sunwoo2023LayerWise, jayaram2022adaptive], adaptive optimizers [reddi2020FedOPT], adaptive Network Layering [karimi2021layer] and adaptive client sampling [Boxin2023AddressingBudget], etc. Wang et al[WangS2019ResourceConstrained] improved the federated learning algorithm by considering scenarios where clients and the server consume the same type of resources. They achieved faster convergence of the global model under resource constraints by dynamically adjusting local iterations.